OpenStudy (anonymous):

multiply (2+5i)(5-2i)

OpenStudy (anonymous):

21i

OpenStudy (anonymous):

The answer is 20+21i.

OpenStudy (anonymous):

Rate if u like:)

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

\[2 \times 5 - 5 \times (-2)=20\] \[2 times (-2)+5 \times 5=21\]

OpenStudy (anonymous):

so answer is 20 + 20i as agarshna said.

OpenStudy (anonymous):

sorry calculation mistake its 20+21i

OpenStudy (radar):

Multiply through getting 10 +25i-4i-10i^2 Consolidate similar terms: 10+21i-10i^2 Take care of the last term as i^2 us equal to a -1 so the last term is 10 final answer is 21i+20

OpenStudy (radar):

or placing the real term first: 20+21i

OpenStudy (radar):

Any questions regarding the procedure for obtaining the solution?